Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability - class of children

Expert replies
by Brent@GMATPrepNow » Sat Jan 31, 2009 2:48 pm
A certain class has an equal number of boys and girls. If three children are chosen at random, the probability that all three children are the same GMAT is 1/5. How many children are in the class?

A) 16
B) 18
C) 22
D) 24
E) 26

Please note that this is not an official GMAT question; it’s my attempt to create difficult (650+ level) GMAT-style questions for this forum.
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion
Source: — Problem Solving |

Re: Probability - class of children

by logitech » Sat Jan 31, 2009 3:08 pm
Boys= X
Girls = X

the probability that all three children are girls = 1/2 x 1/5 = 1/10

C(X:3)/C(2x:3) = 1/10

[x(x-1)(x-2)]/[(2x)(2x-1)(2x-2)] = 1/10

(x-2)/(8x-4) = 1/10

10X-20=8x-4

2x = 16
LGTCH
---------------------
"DON'T LET ANYONE STEAL YOUR DREAM!"
Join the discussion

logitech wrote:Boys= X
Girls = X

the probability that all three children are girls = 1/2 x 1/5 = 1/10

C(X:3)/C(2x:3) = 1/10

[x(x-1)(x-2)]/[(2x)(2x-1)(2x-2)] = 1/10

(x-2)/(8x-4) = 1/10

10X-20=8x-4

2x = 16
Nice work, Logitech

Here's my soln:

Let k be the number of boys. This means that there are k girls and there are 2k children altogether.
Rather than work with both possible cases (all three are boys and all three are girls), it might be easiest to work with one case, say all 3 are girls.

If P(all 3 are same GMAT)=1/5, then P(all three are boys) + P(all three are girls) = 1/5
Since P(all three are boys) = P(all three are girls), then P(all three are girls) = 1/10
P(3 girls are selected) = P(1st child is a girl AND 2nd child is a girl AND 3rd child is a girl)
P(3 girls are selected) = P(1st child is a girl) x P(2nd child is a girl) x P(3rd child is a girl)
P(3 girls are selected) = k/2k x (k-1)/(2k-1) x (k-2)/(2k-2)
P(3 girls are selected) = (k-2)/(8k-4) (simplify)
We are told that P(3 girls are selected) = 1/10
So, we get the equation 1/10 = (k-2)/(8k-4)
Solve for k to get k=8
So, there are 16 children (answer choice A)
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by gaggleofgirls » Sat Jan 31, 2009 7:33 pm
I approached it differently:

X = Total number of children.

The probability that the first child chosen is of one GMAT = x/x = 1

The probability that the second child chosen is the same GMAT as the first = (x/2 -1)/x Because x/2 is all the chhildren of one GMAT and you already used one of them in the first choice.

The probability that the third child chosen matches the same GMAT as the first two = (x/2 -2)/x

Multiply the probabilities together (you can drop the first one, since it = 1) =
(x/2-1)(x/2-2) / (x-1)(x-2) = 1/5

At this point, seems easier just to plug in the answers to try than to work out the math...

(16/2 -1)(16/2-2) / (16-1)(16-2) = 7 * 6 / 15 * 14 = 7 * 3 * 2 / 15 * 7 * 2 = 3/15 = 1/5

Answer = A

-Carrie
Join the discussion

by raptor84 » Sat Jan 31, 2009 9:07 pm
backsolving....

let number of children be 16.

Boys = Girls = 8

Now number of ways 3 children can be selected is 3C16

= 16*15*14/3*2*1 = 8*5*14

Now the number of ways 3 boys can be selected from the 8 boys is

= 8*7*6/3*2*1 = 8*7

Therefore the probability that 3 children selected from the total would all be boys

= 8*7/8*5*14 = 1/10

Similarly solving for number of girls we get = 1/10

Total probability = 1/10+1/10 = 1/5 as provided in the question stem.
Join the discussion